# Improving Sequential Market Clearing via Value-oriented Renewable Energy Forecasting

### v1

@@ Abstract
Large penetration of renewable energy sources (RESs) brings huge uncertainty into the electricity markets. While existing deterministic market clearing fails to accommodate the uncertainty, the recently proposed stochastic market clearing struggles to achieve desirable market properties. In this work, we propose a value-oriented forecasting approach, which tactically determines the RESs generation that enters the day-ahead market. With such a forecast, the existing deterministic market clearing framework can be maintained, and the day-ahead and real-time overall operation cost is reduced. At the training phase, the forecast model parameters are estimated to minimize expected day-ahead and real-time overall operation costs, instead of minimizing forecast errors in a statistical sense. Theoretically, we derive the exact form of the loss function for training the forecast model that aligns with such a goal. For market clearing modeled by linear programs, this loss function is a piecewise linear function. Additionally, we derive the analytical gradient of the loss function with respect to the forecast, which inspires an efficient training strategy. A numerical study shows our forecasts can bring significant benefits of the overall cost reduction to deterministic market clearing, compared to quality-oriented forecasting approach.
@@ I Introduction
A significant challenge with sequential deterministic market clearing arises from its limited ability to address the uncertainty associated with RESs. The separation of DA and RT market clearings means that DA market clearing does not adequately consider the re-dispatch costs incurred due to RES uncertainty [3]. This often results in higher overall operation costs. Consequently, stochastic market clearing is proposed, which aims at informing the DA market with the operation cost in the RT market to reduce overall costs [4]. Though economic efficiency can be improved, stochastic market clearing faces a challenge in simultaneously achieving important market properties, namely revenue adequacy and cost recovery [5]. Therefore, attempts have been made to uphold desirable market properties in stochastic market clearing. For example, [6] ensures cost recovery and revenue adequacy per scenario and in expectation, albeit at the expense of market efficiency.
Alternatively, there has been a growing body of work focusing on tactically scheduling RESs within the current deterministic operation pipeline to emulate the performance of their stochastic program-based counterparts. The core idea is to retain a deterministic DA market clearing model but to tactically schedule RESs by accounting for the future outcomes in RT markets [7, 8]. For example, studies [9, 10] maintain the current deterministic clearing framework and strategically determine wind power dispatch by solving a bilevel program on a case-by-case basis during the operational phase. The study [11] utilizes these methods to allocate reserves according to predicted deviations in RESs. While these methods show a decrease in operation costs, they might pose computational challenges in determining the RES schedule.
This motivates the exploration of the following technical question: Is it possible to train a forecast model (a function, not a fixed solution) that maps the context (e.g. features for RES forecast) to an appropriate RES dispatch schedule, allowing it to enter the deterministic DA market and minimize the overall DA and RT operation costs? In this way, the RES schedule can be conveniently issued by the forecast model during the operational phase. Aligning the training objective of a forecast model with the value of the operation criteria poses a challenge, and is encompassed within the realm of value-oriented forecasting [12, 13, 14].
@@
Several research threads have emerged to address this challenge, encompassing integrated optimization, differentiable programming, and the loss function design. In the first thread, forecast model parameters are optimized concurrently with decision variables. This integrated program is readily solved using commercial solvers [13]. Another approach, proposed by [14], introduces a bilevel program. In this setup, DA operations form the lower level with the forecast as a parameter, while the upper level optimizes the model parameters alongside RT decisions. Notably, this method requires a linear forecast model, potentially limiting its performance. The second thread accommodates more sophisticated forecast models, such as neural networks, by obtaining the gradient of the optimal decision-making objective with respect to the forecast [15]. However, obtaining this gradient involves solving the inverse of Karush-Kuhn-Tucker (KKT) conditions, making it computationally expensive. The third thread, which is the primary focus of our work, centers around the design of loss functions. Existing loss functions, such as pinball loss [16] or SPO loss [17], are primarily tailored for single-stage stochastic programs without redispatch decisions (such as those related to flexible units in the RT market). The design of a value-oriented loss function for sequential market clearing remains an open question.
In this study, our focus lies in the analytical formulation of a loss function specifically crafted for training a value-oriented RES forecast model, aligning with the objective of sequential market clearing. We formulate the parameter estimation task as a bilevel program, optimizing the forecast model parameter at the upper level and the operation decisions are determined at the lower level. Specifically, the DA and RT market clearing problems are solved at the lower level, with the RES forecast from the upper level serving as the input. To show the relationship between the operation cost and the RES forecast more clearly, we resort to the lower-level dual problems and replace the upper-level objective with the dual objectives. To obtain the loss function for training forecast models (whose input is the forecast, and output is the overall DA and RT operation cost), we transform the reformulated upper-level objective as an analytical function of the forecast. Given that the DA and RT market clearings are more general compared to the operational problems analyzed in [18], the reformulated upper-level objective integrates not only the forecast but also the primal and dual solutions derived from the lower level. Hence, it is necessary to derive the functions linking the primal and dual solutions with the forecast. Concretely, we derive the function between dual solutions and forecasts by solving lower-level dual problems. The function between primal solutions and forecasts is derived via the active constraints of the lower-level primal problems. By substituting the primal and dual solutions with these functions, the upper-level objective transforms into a function of the forecast, serving as the loss function for training. Our main contributions are,
@@
1) From a market perspective, the proposed approach maintains the deterministic market clearing framework, while minimizing the overall DA and RT operation cost with value-oriented RES forecasts.
@@
2) From a theoretical perspective, we analytically derive a value-oriented loss function that aligns the training objective of a forecast model with the operation value, i.e., minimizing the overall operation cost.
@@
3) From a practitioner perspective, the analytical loss function allows the analytical derivation of the gradient, which is computationally cheap compared to [15] and inspires a computationally efficient training approach.
@@
The remaining parts of this paper are organized as follows. The preliminaries regarding the sequential market clearing are given in Section II. Section III formulates a bilevel program for forecast model parameter estimation. Section IV derives the loss function for value-oriented forecasting and the training process is presented in Section V. Results are discussed and evaluated in Section VI, followed by the conclusions.
@@ II-A Sequential Market Clearing
We consider the sequential clearing of DA and RT markets [1], which is illustrated in Fig. 1. The DA market is cleared at time $t$ on day $d-1$, with an advance of $k$ hours in time to the next day $d$, and covers energy transactions on day $d$, typically on an hourly basis. Since the RES production is uncertain at the time of the DA market, the energy imbalance with respect to the DA production schedule needs to be settled in the RT market. In line with European practices, we do not incorporate binary decisions regarding unit commitment (UC). However, we note that UC is a requisite consideration in the U.S. DA and RT markets. To analyze the market behavior, the relaxed UC problem, where the binary commitment decisions are substituted with continuous ones, is widely used [3, 10]. Our approach remains applicable to the relaxed problem. More details of the two markets can be found in [1].
Figure 1: An illustration of the sequential market clearing.
Concretely, in the DA market, the operator determines the schedules of generators and RES to satisfy inelastic demand. The generation and RES schedule for each time-slot $\tau,\forall\tau=1,...,24$ on the next day $d$ are denoted as $\bm{p}_{d,\tau}$ and $\bm{w}_{d,\tau}$, respectively. The DA market clearing is formulated as,
$\mathop{\min}_{\Xi_{\text{DA}}}$ $\sum_{\tau=1}^{T}\bm{\rho}^{\top}\bm{p}_{d,\tau}$ (1a)
@@
where $\Xi_{\text{DA}}=\{\bm{p}_{d,\tau},\bm{w}_{d,\tau}\}_{\tau=1}^{T}$. $\bm{\rho}$ is the marginal cost vector of traditional generators. RES enters the market with zero marginal cost. Each element in the vector $\bm{p}_{d,\tau}$ represents the power generated by a generator unit, whose marginal cost is in the corresponding element in $\bm{\rho}$. In the case where generators submit stepwise marginal cost curves (which is an approximation of the linear marginal cost curve [19]), the elements in $\bm{p}_{d,\tau}$ represent different generation blocks with different costs. Here, we assume there is no power loss on lines, and include a DC representation of the network. The equality constraint (1b) enforces the power balance conditions. For simplification, the demand $\bm{l}_{d,\tau}$ is considered to be known with certainty. But we note that this simplification can be easily removed. In this way, the net demand (which is demand minus RES) is uncertain and required to be forecasted. The inequality constraints (1c) restrict the scheduled power flow within the line flow limits. $\bm{H}$ in (1c) is the shift factor matrix mapping the nodal power injection to the power flow on lines [20]. (1d) and (1e) are the output power and ramping limits of the generators. (1f) limits the DA schedule of RES up to the forecast $\hat{\bm{y}}_{d,\tau}$ representing a single-value estimate of the RES production $Y_{d,\tau}$. $Y_{d,\tau}$ is a random variable, since the RES production is unknown at the moment of DA market clearing. After solving (1), the optimal solutions are obtained and denoted as $\{\bm{p}^{*}_{d,\tau},\bm{w}^{*}_{d,\tau}\}_{\tau=1}^{T}$.
@@
where $\Xi_{d,\tau}^{\text{RT}}=\{\bm{p}_{d,\tau}^{+},\bm{p}_{d,\tau}^{-},\bm{\kappa}_{d,\tau}\}$. The output power of generators may be increased by an amount $\bm{p}_{d,\tau}^{+}$ with the marginal cost $\bm{\rho}_{+}$ for up-regulation, or decreased by an amount $\bm{p}_{d,\tau}^{-}$ with the marginal cost $\bm{\rho}_{-}$ for down-regulation. These decisions are driven by the need to settle the RES deviation $\bm{y}_{d,\tau}-\bm{w}_{d,\tau}^{*}$ in (2b). (2c) is the power flow constraint, whose lower and upper bounds are determined by subtracting the power flow in the DA market from the line capacity. (2d) and (2e) limit the amount of up-regulation and down-regulation power to $\overline{\bm{p}^{+}},\overline{\bm{p}^{-}}$. For inflexible generators that cannot be dispatched in RT, the corresponding elements in the upper bounds will be zero, resulting in zero up- and down-adjustments for those generators. Additionally, the eventual generation power, considering the DA schedule $\bm{p}_{d,\tau}^{*}$ and the adjustment, should be within the output power limits, as stated in (2f). The inclusion of RES spill accounts for situations where the actual RES generation surpasses the scheduled amount in the DA schedule $\bm{w}_{d,\tau}^{*}$, and the excess cannot be entirely offset by the down-regulation power available from flexible generators. The amount of RES spill $\bm{\kappa}_{d,\tau}$ can be at most to its realization $\bm{y}_{d,\tau}$, as stated in (2g).
@@
$\bm{\rho}_{+}-\bm{\rho}$ and $\bm{\rho}-\bm{\rho}_{-}$ are the incremental bidding price, which reflects the marginal opportunity loss for up- and down-regulation [5]. We require them to be positive. In this way, the overall generation cost is the minimum if the generators can be dispatched to the eventual schedule in DA. Any RT adjustment would bring the extra cost either $(\bm{\rho}_{+}-\bm{\rho})^{\top}\bm{p}^{+*}_{d,\tau}$ or $(\bm{\rho}-\bm{\rho}_{-})^{\top}\bm{p}^{-*}_{d,\tau}$. The incremental bidding prices of supplying upward and downward balancing power are usually different. This explains why forecasting the expectation hardly works well in reducing the overall cost, as it overlooks the typical asymmetry affecting the RT cost.
@@ II-B Mathematical Reformulation
where $\bm{x}_{d}=[\bm{x}_{d,\tau}]_{\tau=1}^{T}=[\bm{p}_{d,\tau};\bm{w}_{d,\tau}]_{\tau=1}^{T}$ is the collection of DA decision variables. Its optimal solution is denoted as $\bm{x}_{d}^{*}=[\bm{x}_{d,\tau}^{*}]_{\tau=1}^{T}=[\bm{p}_{d,\tau}^{*};\bm{w}_{d,\tau}^{*}]_{\tau=1}^{T}$. The coefficients $\bm{\rho}_{\text{DA}},\bm{G}_{\text{DA}},\bm{\psi}_{\text{DA},d},\bm{F}_{\text{DA}}^{y}$ are constant, whose specific forms are provided in Appendix A. The RES forecasts and the demand are summarized into vectors $\hat{\bm{y}}_{d}=[\hat{\bm{y}}_{d,\tau}]_{\tau=1}^{T}$ and $\bm{l}_{d}=[\bm{l}_{d,\tau}]_{\tau=1}^{T}$, respectively. The value of $\bm{\psi}_{\text{DA},d}$ varies from day to day due to its dependence on the demand $\bm{l}_{d}$. Likewise, the RT market clearing in (7) and (9) are converted into the compact forms,
@@
s.t. $\quad\bm{G}_{\text{RT}}\bm{z}_{d,\tau}\leq\bm{\psi}_{\text{RT},d,\tau}+\bm{F}_{\text{RT}}\bm{x}_{d,\tau}^{*}:$
$\quad\bm{\nu}_{d,\tau},\forall\tau=1.$ (11b)
@@
where $\bm{z}_{d,\tau}=[\bm{p}_{d,\tau}^{+};\bm{p}_{d,\tau}^{-};\bm{\kappa}_{d,\tau}]$ is the collection of RT decision variables. The coefficients $\bm{\rho}_{\text{RT}},\bm{G}_{\text{RT}},\bm{\psi}_{\text{RT},d,\tau},\bm{F}_{\text{RT}}$ and $\bm{G}^{\prime}_{\text{RT}},\bm{\psi}^{\prime}_{\text{RT},d,\tau},\bm{F}^{\prime x}_{\text{RT}},\bm{F}^{\prime p}_{\text{RT}},\bm{F}^{\prime+-}_{\text{RT}}$ are constant and provided in the Appendix A. The values of $\bm{\psi}_{\text{RT},d,\tau},\bm{\psi}_{\text{RT},d,\tau}^{\prime}$ vary from hour to hour due to the dependence on the RES realization $\bm{y}_{d,\tau}$. The parameter $\bm{p}^{+-*}_{d,\tau-1}=[\bm{p}^{+*}_{d,\tau-1};\bm{p}^{-*}_{d,\tau-1}]$ is the collection of RT power adjustment for up- and down- regulation at previous time $\tau-1$.

### v2

# Improving Sequential Market Coordination via Value-oriented Renewable Energy Forecasting

@@ Abstract
Large penetration of renewable energy sources (RESs) brings huge uncertainty into the electricity markets. The current deterministic clearing approach in the day-ahead (DA) market, where RESs participate based on expected production, has been criticized for causing a lack of coordination between the DA and real-time (RT) markets, leading to high overall operating costs. Previous works indicate that improving day-ahead RES entering quantities can significantly mitigate the drawbacks of deterministic clearing. In this work, we propose using a trained forecasting model, referred to as value-oriented forecasting, to determine RES Improved Entering Quantities (RIEQ) more efficiently during the operational phase. Unlike traditional models that minimize statistical forecasting errors, our approach trains model parameters to minimize the expected overall operating costs across both DA and RT markets. We derive the exact form of the loss function used for training, which becomes piecewise linear when market clearing is modeled by linear programs. Additionally, we provide the analytical gradient of the loss function with respect to the forecast, enabling an efficient training strategy. Numerical studies demonstrate that our forecasts significantly reduce overall operating costs for deterministic market clearing compared to conventional forecasts based on expected RES production.
@@ I Introduction
A significant challenge with sequential deterministic market clearing arises from its limited ability to coordinate the two markets under large RES uncertainty. Currently, the DA market is cleared without accounting for its impact on RT market clearing. As a result, DA and RT markets operate without coordination, leading to high overall operating costs [3]. Consequently, stochastic market clearing has been proposed, which informs the DA market of the operating cost in the RT market [4]. While stochastic market clearing can improve overall economic efficiency, it struggles to simultaneously maintain key market properties, such as revenue adequacy and cost recovery [5]. To address this, efforts have been made to preserve these desirable properties within stochastic market clearing. For instance, [6] guarantees cost recovery and revenue adequacy both per scenario and in expectation, though this comes at the cost of reduced market efficiency.
Previous studies have shown that combining DA deterministic market clearing with RES improved entering quantities (RIEQ) enhances coordination between DA and RT markets, achieving performance comparable to stochastic clearing [7, 8]. The key idea is to tactically determine RIEQ—used as input parameters in DA clearing—by considering their impact on future RT market operations. For instance, studies [7, 8] optimize RIEQ by solving a bilevel program on a case-by-case basis during the operational phase. Similarly, [9] applies these methods to predict reserve requirements, facilitating coordination of reserve and energy markets. While these approaches hold promise for improving market coordination, they introduce computational challenges in determining RIEQ.
Building on this foundation, the goal of this work is not to replace the existing RES entering quantities determined by RES producer offerings in the DA market. Instead, we aim to explore an important question: What if the RES entering quantities in the DA market were improved as RIEQ by a “properly” trained forecasting model, specifically optimized to minimize overall system operating costs? If so, could the sequential markets be better coordinated? Here, the overall system operating cost in both the DA and RT markets is used as the quantitative metric to evaluate market coordination.
With the trained forecasting model for determining RIEQ, the computational burden at the operational phase [7, 8] can be avoided. However, aligning the training objective of a forecasting model with the operational goal is a significant challenge, falling within the realm of value-oriented forecasting [10, 11, 12]. Several research threads have emerged to address this challenge, encompassing integrated optimization, differentiable programming, and the loss function design.
@@
For the first thread of research on integrated optimization, forecasting model parameters are optimized concurrently with decision variables [11, 13]. In the same vein, [12] introduces a bilevel program where DA market clearing forms the lower level with the RIEQ prediction as a parameter, while the upper level optimizes both model parameters and RT decisions. This method requires a linear forecasting model to make the program solvable by commercial optimization solvers, which may limit the performance. The second thread of research, on differentiable programming, enables the use of more advanced forecasting models, such as neural networks, by deriving the gradient of decision solution w.r.t. the forecast [14, 15]. However, obtaining such a gradient involves repeatedly solving the inverse of a large-size matrix resulting from the Karush–Kuhn–Tucker conditions, making it computationally expensive. The third approach, and the focus of our work, centers on designing value-oriented loss functions. The existing ones, such as pinball loss [16] or Smart “Predict, then Optimize” (SPO) loss [17], are primarily designed for single-stage stochastic programs (e.g., the one without considering flexible unit redispatch in the RT market). Thus, how to develop a value-oriented loss function for sequential market clearing remains an open question.
To address the challenge of training the RIEQ forecasting model, we propose a tailored loss function that uses the RIEQ forecast as the input and the overall DA and RT system operating costs as the output. The forecasting model parameter estimation is formulated as a bilevel program: the upper level optimizes forecasting model parameters, while the lower level solves the DA and RT market clearing problems using the RIEQ forecast as the input parameter. To link operating cost with the RIEQ forecast, we leverage the lower-level dual problems and replace the upper-level objective with dual objectives. The reformulated upper-level objective is then transformed into an analytical function of the forecast, serving as the loss function for training. This requires the derivation of functions that connect the forecast with primal and dual solutions. Specifically, the dual solutions are linked to forecasts through the dual problems of DA and RT market clearing, while the primal solutions are derived from the active constraints of their respective primal problems. Substituting these functions into the upper-level objective yields a value-oriented loss function for training. Our main contributions are,
@@
1) From a market perspective, the proposed approach maintains deterministic DA market clearing while improving sequential market coordination with RIEQ forecasts.
@@
2) From a theoretical perspective, we derive a value-oriented loss function that aligns the forecasting model’s training objective with operational value, minimizing overall system operating costs in both DA and RT markets.
@@
3) From a practitioner’s perspective, the analytical loss function allows the analytical derivation of the gradient. Leveraging the structure of the gradient, we propose a computationally efficient training approach.
Although prior works [18, 19] design forecasting products using value-oriented loss functions, these approaches are not directly applicable to electricity market settings. First, prior works cannot address the participation of entities other than RES in both DA and RT markets, which is typical in electricity markets. As a result, the network and the redispatch of flexible resources cannot be handled. Second, the generalized market setting introduces new technical challenges in obtaining the value-oriented loss function. The reformulated upper-level objective, derived from dual problems, depends on both the forecast and the primal and dual solutions from the lower level. This requires us to identify the functions between those solutions and the forecast.
@@
The remaining parts of this paper are organized as follows. The preliminaries regarding the sequential market clearing are given in Section II. Section III formulates a bilevel program for forecasting model parameter estimation. Section IV derives the loss function for value-oriented forecasting and the training process is presented in Section V. Results are discussed and evaluated in Section VI, followed by the conclusions.
Notations: The notation $\bm{X}[\mathcal{J}]$ signifies the sub-matrix comprised of rows from the matrix $\bm{X}$ whose indices are included in the index set $\mathcal{J}$. Considering the column vectors $\bm{x}_{1}$ and $\bm{x}_{2}$, the expression $\bm{x}=[\bm{x}_{1};\bm{x}_{2}]$ indicates the vertical concatenation of $\bm{x}_{1}$ and $\bm{x}_{2}$, resulting in the formation of $\bm{x}$. The operator $\Pi_{\mathcal{J}}$ is used to extract a segment from the vector, such as retrieving $\bm{x}_{1}$ from $\bm{x}$.
@@ II Preliminaries
Figure 1: An illustration of forecasting and the sequential market clearing where $\hat{\bm{y}}_{d,\tau}$ and $\bm{y}_{d,\tau}$ are forecasts and realization, and $\bm{w}_{d,\tau}^{*}$ is the RES schedule.
@@ II-A Sequential Market Clearing
We consider the sequential clearing of DA and RT markets [1], as illustrated in Fig. 1. The DA market is cleared at time $t$ on day $d-1$, with an advance of $k$ hours in time to the next day $d$, and covers energy transactions on day $d$, typically on an hourly basis. RES production is uncertain during the DA market, where the RES forecast acts as the RIEQ. Due to inevitable forecasting errors, the energy imbalance caused by forecasting errors needs to be settled in the RT market. Concretely, in the DA market, the operator determines the schedules of generators and RES to satisfy inelastic demand. The generation and RES schedule for each time-slot $\tau,\forall\tau=1,...,24$ on the next day $d$ are denoted as $\bm{p}_{d,\tau}$ and $\bm{w}_{d,\tau}$, respectively. The DA market clearing is,
$\mathop{\min}_{\bm{x}_{d}}$ $\sum_{\tau=1}^{T}\bm{\rho}^{\top}\bm{p}_{d,\tau}$ (1a)
@@
where $\bm{x}_{d}=[\bm{x}_{d,\tau}]_{\tau=1}^{T}=[\bm{p}_{d,\tau}\in\mathbb{R}^{N};\bm{w}_{d,\tau}\in\mathbb{R}^{N}]_{\tau=1}^{T}$ is the collection of DA decision variables, and $N$ is the number of nodes in the system. The optimal solution is denoted as $\bm{x}_{d}^{*}=[\bm{x}_{d,\tau}^{*}]_{\tau=1}^{T}=[\bm{p}_{d,\tau}^{*};\bm{w}_{d,\tau}^{*}]_{\tau=1}^{T}$. $\bm{\rho}\in\mathbb{R}^{N}$ is the marginal cost vector of traditional generators. RES enters the market with zero marginal cost. Each element in the vector $\bm{p}_{d,\tau}$ represents the power generated by a traditional generator unit, whose marginal cost is in the corresponding element in $\bm{\rho}$. Here, we assume there is no power loss on lines, and include a DC representation of the network. The equality constraint (1b) enforces the power balance conditions. For simplification and considering the high accuracy of demand forecast, the demand $\bm{l}_{d,\tau}\in\mathbb{R}^{N}$ is considered to be known with certainty. The inequality constraints (1c) restrict the scheduled power flow within the line flow limits. $\bm{H}$ in (1c) is the Power Transfer Distribution Factors [20] mapping the nodal power injection to the power flow on lines. (1d) and (1e) are the output power and ramping limits of the traditional generators. (1f) limits the DA schedule of RES up to the forecast $\hat{\bm{y}}_{d,\tau}$ which represents a single-value estimate of the RES production $Y_{d,\tau}$. $Y_{d,\tau}$ is a random variable since the RES production is unknown in DA market clearing.
Remark 1.
In line with European practices [12, 21], we do not incorporate binary decisions regarding unit commitment (UC) in (1). However, we note that UC is a requisite consideration in the U.S. markets. To analyze the market behavior, the relaxed UC problem, where the binary commitment decisions are substituted with continuous ones, is widely used [3, 8]. The relaxed UC is a linear program. A common practice is to train the forecasting model using the relaxed UC formulation [8]. During the operational phase, the generated forecasts are then used as inputs to the UC problem.
@@
where $\bm{z}_{d,\tau}=[\bm{p}_{d,\tau}^{+}\in\mathbb{R}^{N};\bm{p}_{d,\tau}^{-}\in\mathbb{R}^{N};\bm{\kappa}_{d,\tau}\in\mathbb{R}^{N}]$ is the collection of RT decision variables. The output power of generators may be increased by an amount $\bm{p}_{d,\tau}^{+}$ with the marginal cost $\bm{\rho}_{+}>0$ for up-regulation, or decreased by an amount $\bm{p}_{d,\tau}^{-}$ with the marginal utility $\bm{\rho}_{-}>0$ for down-regulation. $\bm{\kappa}_{d,\tau}$ is the amount of RES spill. These decisions are driven by the need to settle the RES deviation $\bm{y}_{d,\tau}-\bm{w}_{d,\tau}^{*}$ in (2b). (2c) is the power flow constraint, whose lower and upper bounds are determined by subtracting the power flow in the DA market from the line capacity. (2d) and (2e) limit the amount of up-regulation and down-regulation power to $\overline{\bm{p}^{+}},\overline{\bm{p}^{-}}$. For inflexible generators that cannot be dispatched in RT, the corresponding elements in the upper bounds will be zero, resulting in zero up- and down-adjustments for those generators. Additionally, the eventual generation power, considering the DA schedule $\bm{p}_{d,\tau}^{*}$ and the adjustment, should be within the output power limits, as stated in (2f). The inclusion of RES spill accounts for situations where the actual RES generation surpasses the scheduled amount in the DA schedule $\bm{w}_{d,\tau}^{*}$, and the excess cannot be entirely offset by the down-regulation power available from flexible generators. The amount of RES spill $\bm{\kappa}_{d,\tau}$ can be at most to its realization $\bm{y}_{d,\tau}$, as stated in (2g).
@@
$\bm{\rho}_{+}-\bm{\rho}$ and $\bm{\rho}-\bm{\rho}_{-}$ are the incremental bidding price, which reflects the marginal opportunity loss for up- and down-regulation [5]. We require them to be positive for penalizing the RT adjustment due to forecasting errors. In this way, any RT adjustment would bring the extra cost either $(\bm{\rho}_{+}-\bm{\rho})^{\top}\bm{p}^{+*}_{d,\tau}$ or $(\bm{\rho}-\bm{\rho}_{-})^{\top}\bm{p}^{-*}_{d,\tau}$. The incremental bidding prices of supplying upward and downward balancing power are usually different. This explains why forecasting the expectation hardly works well in reducing the overall cost, as it overlooks the typical asymmetry affecting the RT cost.
@@ II-B Mathematical Reformulation
The coefficients $\bm{\rho}_{\text{DA}},\bm{G}_{\text{DA}},\bm{\psi}_{\text{DA},d},\bm{F}_{\text{DA}}^{y}$ are constant. The RES forecasts and the demand are summarized into vectors $\hat{\bm{y}}_{d}=[\hat{\bm{y}}_{d,\tau}]_{\tau=1}^{T}$ and $\bm{l}_{d}=[\bm{l}_{d,\tau}]_{\tau=1}^{T}$, respectively. The value of $\bm{\psi}_{\text{DA},d}$ varies from day to day due to its dependence on the demand $\bm{l}_{d}$. Likewise, the RT market clearing in (7) and (9) are converted into the compact forms,
@@
s.t. $\bm{G}_{\text{RT}}\bm{z}_{d,\tau}\leq\bm{\psi}_{\text{RT},d,\tau}+\bm{F}_{\text{RT}}\bm{x}_{d,\tau}^{*}:\bm{\nu}_{d,\tau},\tau=1.$ (11b)
@@
The coefficients $\bm{\rho}_{\text{RT}},\bm{G}_{\text{RT}},\bm{\psi}_{\text{RT},d,\tau},\bm{F}_{\text{RT}}$ and $\bm{G}^{\prime}_{\text{RT}},\bm{\psi}^{\prime}_{\text{RT},d,\tau},\bm{F}^{\prime x}_{\text{RT}},\bm{F}^{\prime p}_{\text{RT}},\bm{F}^{\prime+-}_{\text{RT}}$ are constant. The values of $\bm{\psi}_{\text{RT},d,\tau},\bm{\psi}_{\text{RT},d,\tau}^{\prime}$ vary from hour to hour due to the dependence on the RES realization $\bm{y}_{d,\tau}$. The parameter $\bm{p}^{+-*}_{d,\tau-1}=[\bm{p}^{+*}_{d,\tau-1};\bm{p}^{-*}_{d,\tau-1}]$ contains RT power adjustment for up- and down- regulation at previous time $\tau-1$.